2. Phosphorus (V) oxide P4O10 is used to prepare phosphoric acid.

a) Write the reaction equation when solid phosphorus (V) oxide is prepared by burning solid phosphorus P4.

b) How much phosphorus (V) oxide can be produced when 1.33 g of phosphorus and 5.07 g of oxygen are available?

3.
Butane with an evaporation temperature of 0°C is used as liquefied gas in camping stoves. Butane is also a great power for summer hikers in Finland as a gas for a camping stove. One summer day the temperature was + 15°C and the air pressure 0.95 bar. Calculate how many liters of carbon dioxide was formed when a hiker boiled a 450-gram bottle of butane gas. Suppose the combustion was complete and that the hiker managed to cook the bottle empty (in fact, the same pressure remained inside the bottle as it was outside the bottle).
1
Expert's answer
2020-12-28T04:41:04-0500

2.

a) P4 + 5O2 = 2P2O5

b) M (P4) = 123.89 g/mol

M (O2) = 32 g/mol

M (P2O5) = 283.89 g/mol

n (P4)= 1.33/123.89 = 0.01 mol

n (O2) = 5.07/32 = 0.16 mol

P4 is the limiting reagent in this reaction.

n (P2O5) = 2 x n (P4) = 2 x 0.01 = 0.02 mol

m (P2O5) = n x M = 0.02 x 283.89 = 5.68 g


3.

2C4H10 + 13O2= 8CO2 + 10H2O

M (C4H10) = 58.12 g/mol

T = 15 + 273 = 288 K

n (C4H10) = 450/58.12 = 7.74 g/mol

n (CO2) = 8/2 n (C4H10) = 8/2 x 7.74 = 30.96 mol

pV=nRT

R = 8.31 L⋅bar⋅K−1⋅mol−1

T = 15 + 273 = 288 K

V (CO2) = nRT/p = (30.96 x 3.14 x 288) / 0.95 = 29 471 L